OpenStudy (anonymous):

In an experiment 1.0912 g of ferric oxide is reacted with 0.4373 g of carbon monoxide gas according to the equation: Fe2O3 + CO ==> Fe + CO2. A. What is the limiting reactant? B. What mass of the other reactant is in exess? C. What mass of iron is produced?

OpenStudy (sweetburger):

Well first you need to balance the equation.

OpenStudy (anonymous):

please can you help me balance the equation

OpenStudy (somy):

look at number of each atom on left side and on right side try to make it so number of atoms on right side is equal to left side

OpenStudy (anonymous):

help me this is my first time of doing chemtry

OpenStudy (somy):

well look Fe2O3 + CO ==> Fe + CO2. lets take O look at left of the arrow tell me how many O atoms do u see?

OpenStudy (somy):

u wrote same thing that was there balance ittt

OpenStudy (anonymous):

i see two o

OpenStudy (somy):

why? i said LEFT where on left side u see 2?

OpenStudy (somy):

why are you jumping <.< where is the other part of the reaction?

OpenStudy (somy):

you did it on your own?

OpenStudy (somy):

well done :) yes that is right :)

OpenStudy (somy):

find mole of Fe2O3 and mole of CO their masses are given

OpenStudy (somy):

mole= mass/ molecular mass

OpenStudy (anonymous):

is this correct 1/ 291.7167 g/mol

OpenStudy (somy):

why 1 divided by that? i suggest you to write the whole mass that they gave, i mean with all decimals

OpenStudy (anonymous):

is that the right answer 291.7167 g/mol

OpenStudy (somy):

i don't have periodic table near me u do know how to take molecular mass right? if so then yeah you are right probably

OpenStudy (somy):

sorry gtg @sweetburger will help out with the rest :)

OpenStudy (anonymous):

ok thank u so much for helping this is my first time of being here how do i tell sweetburger to help me

OpenStudy (anonymous):

are u with me

OpenStudy (somy):

i'll just quickly tell the whole work A- find moles of both reactants => Compound with smallest mole is Limiting reagent B- here you just do some math comparison stuff C- you take the mole of the limiting reagent and check the ratio between limiting reagent and Fe (meaning look at coefficients) and the use this ratio to find mole of Fe from mole using formula mass=mole* Molecular mass get the mass of Fe @sweetburger will explain these steps in depth good luck ❤

OpenStudy (anonymous):

thank you so much may god bless you

OpenStudy (somy):

hehe no probs :)))

OpenStudy (somy):

and thank you ❤

OpenStudy (anonymous):

bye have a wounderfull day

OpenStudy (somy):

night rather lol :))

OpenStudy (anonymous):

someone help

OpenStudy (abb0t):

Did you balance this yet?

OpenStudy (abmon98):

Step 1: Balance Your Chemical Equation Write down the number of atoms that you have on each side of the equation. Look at the subscripts next to each atom to find the number of atoms in the equation. Fe:2 Fe:1 O:4 O:2 C:1 C:1 Start with Oxygen, Carbon and Iron. Carbon Monoxide is a reducing, it would take out all the oxygen atoms from iron. *2 to the Oxygen in CO2 of the product side Fe2O3(s) + CO(g) --> Fe(s) + 2CO2(g) *2 to CO on the reactant side to balance Fe2O3(s) + 2CO(g) --> Fe(s) + 2CO2 Change the Oxygen by *3 instead of 2 Fe2O3(s) + 2CO(g) --> Fe(s) + 3CO2 Change the Oxygen and Carbon on the reactant side Fe2O3(s) + 3CO(g) --> 2Fe(s) + 3CO2 Balance Iron on both side by *2 on reactant side and product side. Fe2O3(s) + 3CO(g) --> 2Fe(s) + 3CO2(g) Number of Moles=Mass(g)/Molar Mass(g/mol) Molar Mass:(55.8*2)+(16*3)=159.6 1.0912/159.6=0.00687 moles of Fe2O3 Molar Mass: 0.4373/(12+16)=0.0156 moles of CO Look at the Mole Ratio of Fe2O3 to CO based on CO mass 1:3 0.00687:x x=0.02061 mole of CO, but there is 0.0156 Look At the Mole Ratio of Fe2O3 to CO based on Fe2O3 1:3 x:0.0156 x=0.00299 This is less thats whats required to react and therefore CO is the limiting reagent. The excess reactant is FeO3 0.00687-0.00299=0.00388 Moles are in Excess 0.00388*159.6=Mass Mass:0.61692 grams of iron Lets get to the last part of your question Look At the Mole Ratio of CO:Fe 3:2 0.0156:x X:0.0156*2/3=0.0104 Moles of iron Are produced. Number of moles=Mass(g)/Molar Mass(g/mol) 0.0104*55.8=Mass(g)

OpenStudy (somy):

The excess reactant is Fe2O3* Well done @Abmon98 :DDD

OpenStudy (abmon98):

Thank you @Somy and sorry for that :/

OpenStudy (anonymous):

please which one is the right answer

OpenStudy (somy):

:)))